OpenStudy (anonymous):

Please help! 3(cos90° + i sin90°) dot (cos39°+ i sin39°)

OpenStudy (anonymous):

i got to the step: 3(cos181° + i sin78°) I'm not sure if i should just add 181° and 78° together ?

OpenStudy (anonymous):

i m getting -3sin39+3icos39

OpenStudy (anonymous):

what?? -.-

OpenStudy (anonymous):

can you please write out the steps ?

OpenStudy (anonymous):

cos90=0 and sin90=1, therefore u get 3(0+i).(cos39+isin39) 3i(cos39+isin39) now since i^2=-1, u get 3icos39+3i^2sin39 -3sin39+3icos39

OpenStudy (anonymous):

oh my god I am soooo sorry, I wrote the problem wrong... it is 3(cos91 + i sin91) not 90 =.=

OpenStudy (anonymous):

my fault so sorry :(

OpenStudy (anonymous):

i was like so weird how the heck u get -3 :D

OpenStudy (anonymous):

now i m getting cos130+isin130

OpenStudy (anonymous):

is my answer correct?

OpenStudy (anonymous):

the heck :D, is that the final answer ?

OpenStudy (anonymous):

i got 3 cis 260° as the answer but I'm not sure :|

OpenStudy (anonymous):

let me recheck my calculations

OpenStudy (anonymous):

ok, but before i got 3 cis 260°. i got 3(cos182°+ i sin 78°) and i just added those two together to get 3 cis 260° but I;m not sure

OpenStudy (anonymous):

can u tell me what does cis mean?

OpenStudy (anonymous):

sorry,forgot 3, my final ans is 3(cos130+isin130)

OpenStudy (anonymous):

(cos91+isin39).(cos39+isin39)=cos91cos39+i(sin91cos39+cos91sin39)-sin39sin91 now cos91cos39-sin91sin39=cos(91+39) and sin91cos39+cos91sin39=sin(91+39) so you get cos130+isin130 multiply the above with 3, this is how i got to my answer

OpenStudy (anonymous):

how did u get to ur ans?

OpenStudy (anonymous):

yeah, got to know what cis means here then my answer is 3cis130 but how did u get 182 and 78 and why did u add them? u can't simply add them.. further cistheta=cos(theta)+isin(theta) since theta is the arguement of the complex number,it has to be the same in both cos and sin terms, u cannot have one value of theta in a particlar term and another in the other term and there is no point in adding them the way u did